# LLMs Are In-Context Bandit Reinforcement Learners

### v2

@@ Acknowledgments
We thank Yair Feldman for proposing Spearman’s rank correlation as a stability metric, and Mustafa Omer Gul, Yair Feldman, Yilun Hua, and Robert West for insightful discussion and feedback.
This research was supported by NSF under grants No. 1750499 and OAC-2311521, NASA under award No. 20-OSTFL20-0053, a gift from Open Philanthropy, a gift from Apple, the National Artificial Intelligence Research Resource (NAIRR) Pilot, the Frontera supercomputer supported by the National Science Foundation (award NSF-OAC 1818253) at the Texas Advanced Computing Center (TACC) at The University of Texas at Austin, and the Delta advanced computing and data resource which is supported by the National Science Foundation (award NSF-OAC 2005572). We thank Google for enabling experiments with Gemini through a gift. AB gratefully acknowledges the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Center for Imaging, Sony Group Corporation, and the Allen Institute for AI. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation, NASA, or the other funders.
@@ B.3.1 Stochastic ICRL Computational Costs
An important technical difference between the Naive and Naive+ approaches and Stochastic is that, until the context window is not saturated, Naive approaches can potentially re-use past computations from caching. This is not possible in Stochastic, because each episode requires the construction of a fresh context $C^{(t)}$. The probability of encountering the same context twice, or even the same prefix, is exceptionally low even after a few episodes. This means that the context has to be computed from scratch for each input. [^10] In low-memory setups, this does not lead to noticeable slowdowns (as efficient caching would not be possible), and Stochastic can be much faster given that each context contains only $p_{\text{keep}}$of the episodes that Naive+ would use. In our setup, we empirically find this to be the case and Stochastic is significantly faster in practice.
@@ Appendix D Additional Results
Table 4: Detailed Figures for Figure 4. We report three key figures for each dataset and method: initial (zero-shot) accuracy, final (post-ICRL) accuracy, and regret (total mistakes). (a) contains the results for Llama, (b) for Qwen, and (c) for Gemini.
@@
Method 0-step Final Reg. 0-step Final Reg. 0-step Final Reg. 0-step Final Reg. 0-step Final Reg.
Acc. Acc. Acc. Acc. Acc. Acc. Acc. Acc. Acc. Acc.
Naive 0.020 0.018 9909 0.002 0.000 9928 0.010 0.028 9674 0.030 0.086 4032 0.034 0.030 4828
Naive+ (w/o Ex.) 0.020 0.454 7164 0.008 0.276 8507 0.010 0.476 6789 0.118 0.790 2046 0.080 0.144 4597
Naive+ (w/ Ex.) 0.024 0.178 4704 0.000 0.590 5727 0.056 0.656 4166 0.238 0.792 2214 0.006 0.196 4153
Stochastic (w/o Ex.) 0.018 0.350 8280 0.002 0.162 9437 0.010 0.428 7288 0.030 0.838 2244 0.034 0.258 4767
Stochastic (w/ Ex.) 0.030 0.584 4380 0.006 0.650 6276 0.162 0.722 3725 0.238 0.840 2424 0.008 0.036 4773
@@
Method 0-step Final Reg. 0-step Final Reg. 0-step Final Reg. 0-step Final Reg. 0-step Final Reg.
Acc. Acc. Acc. Acc. Acc. Acc. Acc. Acc. Acc. Acc.
Naive 0.016 0.002 9877 0.012 0.014 9892 0.010 0.014 9731 0.124 0.13 3876 0.010 0.024 4810
Naive+ (w/o Ex.) 0.014 0.442 7391 0.008 0.204 8824 0.014 0.622 6479 0.136 0.526 3304 0.010 0.146 3713
Naive+ (w/ Ex.) 0.248 0.626 4618 0.048 0.602 4481 0.162 0.656 3818 0.158 0.776 1564 0.004 0.526 3044
Stochastic (w/o Ex.) 0.016 0.208 8653 0.012 0.134 9449 0.010 0.564 6128 0.124 0.886 1591 0.010 0.532 3868
Stochastic (w/ Ex.) 0.316 0.680 4353 0.080 0.712 4385 0.182 0.706 3882 0.200 0.890 1520 0.002 0.462 4440
@@
Method 0-step Final Reg. 0-step Final Reg. 0-step Final Reg. 0-step Final Reg. 0-step Final Reg.
Acc. Acc. Acc. Acc. Acc. Acc. Acc. Acc. Acc. Acc.
Naive 0.018 0.000 9996 0.008 0.010 9929 0.048 0.004 9938 0.076 0.012 4841 0.002 0.018 4915
Naive+ (w/o Ex.) 0.014 0.174 8858 0.002 0.144 9264 0.044 0.218 7978 0.104 0.900 1308 0.006 0.130 3998
Naive+ (w/ Ex.) 0.430 0.778 2776 0.356 0.818 1711 0.474 0.774 2582 0.480 0.896 1127 0.134 0.682 2448
Stochastic (w/o Ex.) 0.014 0.214 8625 0.008 0.060 9710 0.050 0.290 8281 0.072 0.876 1505 0.002 0.060 4920
Stochastic (w/ Ex.) 0.440 0.792 3292 0.434 0.858 1984 0.480 0.794 2560 0.494 0.872 1439 0.138 0.636 3420

### v3

@@ Acknowledgments
We thank Yair Feldman for proposing Spearman’s rank correlation as a stability metric, and Mustafa Omer Gul, Yair Feldman, Yilun Hua, and Robert West for insightful discussion and feedback. This research was supported by NSF under grants No. 1750499 and OAC-2311521, NASA under award No. 20-OSTFL20-0053, a gift from Open Philanthropy, a gift from Apple, the National Artificial Intelligence Research Resource (NAIRR) Pilot, the Frontera supercomputer supported by the National Science Foundation (award NSF-OAC 1818253) at the Texas Advanced Computing Center (TACC) at The University of Texas at Austin, and the Delta advanced computing and data resource which is supported by the National Science Foundation (award NSF-OAC 2005572). We thank Google for enabling experiments with Gemini through a gift. AB gratefully acknowledges the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Center for Imaging, Sony Group Corporation, and the Allen Institute for AI. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation, NASA, or the other funders.
@@ B.3.1 Stochastic ICRL Computational Costs
An important technical difference between the Naive and Naive+ approaches and Stochastic is that, until the context window is not saturated, Naive approaches can potentially re-use past computations from caching. This is not possible in Stochastic, because each episode requires the construction of a fresh context $C^{(t)}$. The probability of encountering the same context twice, or even the same prefix, is exceptionally low even after a few episodes. This means that the context has to be computed from scratch for each input. [^10] In low-memory setups, this does not lead to noticeable slowdowns (as efficient caching would not be possible), and Stochastic can be much faster given that each context contains only $p_{\text{keep}}$% of the episodes that Naive+ would use. In our setup, we empirically find this to be the case and Stochastic is significantly faster in practice.
@@ Appendix D Additional Results
Table 4: Detailed Figures for Figure 4. We report three key figures for each dataset and method: initial (zero-shot) accuracy, final (post-ICRL) accuracy, and regret (total mistakes).
@@
Method 0-step Final 0-step Final 0-step Final 0-step Final 0-step Final
Naive 0.020 0.018 0.002 0.000 0.010 0.028 0.030 0.086 0.034 0.030
Naive+ (w/o Ex.) 0.020 0.454 0.008 0.276 0.010 0.476 0.118 0.790 0.080 0.144
Naive+ (w/ Ex.) 0.024 0.178 0.000 0.590 0.056 0.656 0.238 0.792 0.006 0.196
Stochastic (w/o Ex.) 0.018 0.350 0.002 0.162 0.010 0.428 0.030 0.838 0.034 0.258
Stochastic (w/ Ex.) 0.030 0.584 0.006 0.650 0.162 0.722 0.238 0.840 0.008 0.036
@@
Method Reg. Reg. Reg. Reg. Reg.
Naive 9909 9928 9674 4032 4828
Naive+ (w/o Ex.) 7164 8507 6789 2046 4597
Naive+ (w/ Ex.) 4704 5727 4166 2214 4153
Stochastic (w/o Ex.) 8280 9437 7288 2244 4767
Stochastic (w/ Ex.) 4380 6276 3725 $24$ 4773
Abs. Banking77 Abs. CLINC150 Abs. NLU Abs. TREC Abs. TREC-fine
Method 0-step Final 0-step Final 0-step Final 0-step Final 0-step Final
Naive 0.016 0.002 0.012 0.014 0.010 0.014 0.124 0.130 0.010 0.024
Naive+ (w/o Ex.) 0.014 0.442 0.008 0.204 0.014 0.622 0.136 0.526 0.010 0.146
Naive+ (w/ Ex.) 0.248 0.626 0.048 0.602 0.162 0.656 0.158 0.776 0.004 0.526
Stochastic (w/o Ex.) 0.016 0.208 0.012 0.134 0.010 0.564 0.124 0.886 0.010 0.532
Stochastic (w/ Ex.) 0.316 0.680 0.080 0.712 0.182 0.706 0.200 0.890 0.002 0.462
@@
Method Reg. Reg. Reg. Reg. Reg.
Naive 9877 9892 9731 3876 4810
Naive+ (w/o Ex.) 7391 8824 6479 3304 3713
Naive+ (w/ Ex.) 4618 4481 3818 1564 3044
Stochastic (w/o Ex.) 8653 9449 6128 1591 3868
Stochastic (w/ Ex.) 4353 4385 3882 1520 4440
Abs. Banking77 Abs. CLINC150 Abs. NLU Abs. TREC Abs. TREC-fine
Method 0-step Final 0-step Final 0-step Final 0-step Final 0-step Final
Naive 0.018 0.000 0.008 0.010 0.048 0.004 0.076 0.012 0.002 0.018
Naive+ (w/o Ex.) 0.014 0.174 0.002 0.144 0.044 0.218 0.104 0.900 0.006 0.130
Naive+ (w/ Ex.) 0.430 0.778 0.356 0.818 0.474 0.774 0.480 0.896 0.134 0.682
Stochastic (w/o Ex.) 0.014 0.214 0.008 0.060 0.050 0.290 0.072 0.876 0.002 0.060
Stochastic (w/ Ex.) 0.440 0.792 0.434 0.858 0.480 0.794 0.494 0.872 0.138 0.636
@@
Abs. Banking77 Abs. CLINC150 Abs. NLU Abs. TREC Abs. TREC-fine
Method Reg. Reg. Reg. Reg. Reg.
Naive 9996 9929 9938 4841 4915
Naive+ (w/o Ex.) 8858 9264 7978 1308 3998
Naive+ (w/ Ex.) 2776 1711 2582 1127 2448
Stochastic (w/o Ex.) 8625 9710 8281 1505 4920
Stochastic (w/ Ex.) 3292 1984 2560 1439 3420